How long would n $800 investment take to earn $160 at a rate of 5% annually?



What is the interest rate if you earned $504 on an $1800 investment for 4 years?

Answer by Boreal(15235)

It would then be $960
960=800(1+.05)^n
divide by 800 both sides
1.2=1.05^n
ln 1.2=n ln 1.05
n=ln 1.2/ln 1.05=3.11 years
2304=1800(1+r)^4
1.28=(1+r)^4
ln 1.28=4 ln (1+r)
ln(1+r)=0.0617, don't round
raise everything to the e power
1+r=e^0.0617=1.064
r=0.064 or 6.4%